# Universally Robust Control of Open Quantum Systems

### v1

@@ II.4 Performance Evaluation of Universally Robust Control
To validate our framework, we implement the universally robust control protocol for two fundamental quantum tasks: (i) State transfer and (ii) Quantum gate operations under environmental noise. In the simulations of these control tasks, we model the environment as a thermal bath of harmonic oscillators with $\hat{H}_{\rm B}=\sum_{k}\left(\omega_{k}\hat{a}_{k}^{\dagger}\hat{a}_{k}+\frac{1}{2}\right)$, where $\hat{a}_{k}$ are bosonic annihilation operators for mode $k$ Breuer and Petruccione (2002). The system-environment coupling has the generic form $\hat{H}_{\rm I}=g\hat{A}\otimes\hat{B}$, where $\hat{B}=\sum_{k}\left(g_{k}/g\right)(\hat{a}_{k}+\hat{a}_{k}^{\dagger})$ and a super-Ohmic spectral density $J(\omega)=\omega_{\rm c}^{-2}\omega^{3}{\rm e}^{-\omega/\omega_{\rm c}}$ is assumed Weiss (2021). Here, $\omega_{\rm c}$ is the cutoff frequency, chosen as 10 times of the typical Bohr frequency of the bare system. Such a parameters choice corresponds to a large class of quantum systems Alkauskas et al. (2014); Jahnke et al. (2015); Norambuena et al. (2016). Besides, all calculations employ atomic units (a.u.) in the following.
@@ II.4.1 State Transfer in a Two-Level System
where $\hat{\sigma}_{i}$ ($i=x,y,z$) denote Pauli matrices, $\Delta$ represents the energy splitting, and $u_{x,\,y}(t)$ correspond to independent control fields. This system exhibits complete controllability as the drift Hamiltonian ($\hat{\sigma}_{z}$) and the control Hamiltonians ($\hat{\sigma}_{x,\,y}$) generate the full $\mathfrak{su}(2)$ Lie algebra Huang et al. (1983); Kallush and Kosloff (2011).
@@
We implement quantum state transfer between orthogonal maximally coherent states: initial state $\hat{\rho}_{-x}=\frac{1}{2}(\hat{\mathbb{I}}-\hat{\sigma}_{x})$ [Bloch vector (1,0,0)(-1,0,0)] and target state $\hat{\rho}_{+x}=\frac{1}{2}(\hat{\mathbb{I}}+\hat{\sigma}_{x})$ [Bloch vector $(+1,0,0)$]. To evaluate protocol robustness, we consider two distinct noise channels: (i) specific coupling $\hat{A}=\hat{\sigma}_{z}$ and (ii) generic stochastic coupling $\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$ with $\mathbf{n}$ being a randomly oriented unit vector. Control fields were optimized by minimizing the objective functional $\mathcal{J}$ [Eq. (17)] via the CRAB algorithm, comparing target-only ($c=0$) versus robust ($c>0$) control strategies. Crucially, the identical control fields optimized for $\hat{A}=\hat{\sigma}_{z}$ were applied to the generic coupling case $\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$ without re-optimization to demonstrate universality.
@@
Fig. 1 highlights the striking advantages of our universally robust protocol. For a specific noise channel $\hat{A}=\hat{\sigma}_{z}$ [Fig. 1 (a)], both protocols achieve near-unity fidelity ($\mathcal{F}_{\text{state}}\approx 1$) at zero system-bath coupling ($\lambda=0$). However, under increasing $\lambda$, the target-only optimization exhibits rapid fidelity degradation $\mathcal{F}_{\text{state}}<0.7$ at $\lambda=0.1$], while the robust protocol maintains $\mathcal{F}_{\text{state}}>0.99$ throughout $\lambda\in[0,0.1]$. More importantly, when faced with generic, uncharacterized noise $\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$ [Fig. 1 (b)], the robust control preserves high fidelity ($\mathcal{F}_{\text{state}}>0.99$ at $\lambda=0.1$), whereas the target-only approach fails catastrophically ($\mathcal{F}_{\text{state}}<0.7$). This clearly confirms the universality of our robust control protocol. Beyond its superior fidelity, the robust protocol is also significantly more efficient, requiring control field amplitudes that are approximately 50% lower than the target-only method [Fig. 1 (c)]. Furthermore, it excels at preserving quantum state purity. As illustrated by the Bloch sphere trajectories [Fig. 1 (d)], the robust evolution remains confined to the sphere’s surface, indicating a pure, unitary process, whereas the target-only path spirals inward—a clear signature of decoherence. These combined features demonstrate a capacity for universal noise resistance with minimal operational overhead.
@@
Figure 2: Universally robust control for single-qubit Hadamard gate. Gray dash-dotted lines: target-only control ($c=0$). Red solid lines: universally robust control ($c=10^{-2}$). (a) Gate fidelity $\mathcal{F}_{\rm gate}$ versus coupling strength $\lambda$ for specific noise ($\hat{A}=\hat{\sigma}_{z}$). Inset: Log-log infidelity ($1-\mathcal{F}_{\rm gate}$). (b) $\mathcal{F}_{\rm gate}$ versus $\lambda$ for generic noise ($\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$; $\mathbf{n}$: random unit vector). Data averaged over 20 realizations. Inset: Corresponding infidelity. (c) Optimized control fields $u_{x}(t)$ and $u_{y}(t)$ for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. (d) State evolution trajectories on Bloch sphere under the Hadamard gate (from the z-z to the x-x direction) for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. Initial state: black square. Final states: blue dot (target-only) and green diamond (universally robust control). Parameters: Energy splitting $\Delta=3\times 10^{-3}$ a.u., inverse temperature $\beta=1/\Delta$.
Figure 3: Universally robust control for two-qubit CZ Gate. Gray dash-dotted lines: target-only control ($c=0$). Purple solid lines: universally robust control ($c=2\times 10^{-3}$). (a) Gate fidelity $\mathcal{F}_{\rm gate}$ versus coupling strength $\lambda$ for specific noise ($\hat{A}=\hat{\sigma}_{y}\otimes\hat{\sigma}_{0}$). Inset: Log-log infidelity ($1-\mathcal{F}_{\rm gate}$). (b) $\mathcal{F}_{\rm gate}$ versus $\lambda$ for generic noise $\hat{A}=\sum_{\mu\nu}a_{\mu\nu}\hat{\sigma}_{\mu}\otimes\hat{\sigma}_{\nu}$, where the random coefficients $a_{\mu\nu}\sim N(0,1)$ with $\sum_{\mu\nu}a^{2}_{\mu\nu}=1$]. Data averaged over 20 realizations. Inset: Corresponding infidelity. (c) Optimized control fields $u_{j}(t)$ ($j=1,2,3,4$) for $\hat{A}=\hat{\sigma}_{y}\otimes\hat{\sigma}_{0}$ and $\lambda=0.1$. Parameters: Energy splitting $\Delta=3\times 10^{-3}$ a.u., inverse temperature $\beta=1/\Delta$.
@@ II.4.2 Quantum Gate Operations
We extend our robust control protocol to the more demanding task of quantum gate synthesis, which require noise-resilient state transformations across the entire Hilbert space—a significantly more complex challenge than single-state transfer Palao and Kosloff (2003); Aspman et al. . We demonstrate protocol effectiveness for two fundamental gates: the single-qubit Hadamard gate and two-qubit controlled-Z (CZ) gate, which form a universal gate set when combined with T gates Nielsen and Chuang (2010).
@@
Using the control Hamiltonian from Eq. (18), we observe performance trends mirroring the two-level state transfer results. The influence of the noise strength $\lambda$ on the gate fidelity with $\hat{A}=\hat{\sigma}_{z}$ and $\hat{A}={\bf n}\cdot\hat{\boldsymbol{\sigma}}$ are depicted in Fig. 2 (a) and (b), respectively. In both cases, the target-only optimization results deviate substantially from the ideal value once $\lambda\neq 0$, while those of robust control exhibit strong noise robustness (e.g., preserves $\mathcal{F}_{\text{gate}}>0.99$ at $\lambda=0.1$ [Fig. 2 (b)]). Compared to the target-only optimization, the fidelity degradation of the robust protocol is suppressed by two orders of magnitude. The comparison of the required control fields in the two protocols is shown in Fig. 2 (c), where we find again that the robust optimization requires much lower control-field amplitude than the target-only one. Besides, robust trajectories maintain confinement to the Bloch sphere surface (purity preservation), contrasting with target-only penetration [Fig. 2 (d)].
@@
where $\hat{\sigma}_{0}$ is the identity operator for qubit. The control fields to be optimized are represented by $u_{j}(t)$ ($j=1,2,3,4$). Fig. 3 shows the variation of the gate fidelities against $\lambda$ for (a) $\hat{A}=\hat{\sigma}_{y}\otimes\hat{\sigma}_{0}$ and (b) $\hat{A}=\sum_{\mu\nu}a_{\mu\nu}\hat{\sigma}_{\mu}\otimes\hat{\sigma}_{\nu}$. Here, $\mu,\,\nu=0,x,y,z$, and the random coefficients $a_{\mu\nu}\sim N(0,1)$ with $\sum_{\mu\nu}a^{2}_{\mu\nu}=1$. In both cases, we find qualitatively the same behaviors as those of the single-qubit gate. That is, the results of the target-only optimization decreases rapidly as $\lambda$ increases, while the robust control can achieve and maintain high fidelity across a considerable range of noise amplitudes (e.g., preserves $\mathcal{F}_{\text{gate}}>0.99$ at $\lambda=0.1$ [Fig. 3 (b)]). Besides, as shown in Fig. 3 (c), the robust control requires significantly lowered control-field amplitudes. The robust protocol consistently outperforms target-only optimization across all metrics, demonstrating scalability to multi-qubit systems while maintaining noise resilience and control efficiency.
@@ III Discussion
We have developed and validated a universally robust control framework capable of suppressing noise in open quantum systems without requiring specific knowledge of the noise channels. By engineering a control-objective functional that minimizes an intrinsic noise-susceptibility metric, our method achieves provable robustness against arbitrary Markovian noise. Numerical simulations of state transfer and quantum gate operations show that this approach yields near-unity fidelities and suppresses errors by orders of magnitude compared to target-only optimization, even when subjected to unanticipated noise sources.
This work represents a significant paradigm shift for robust quantum control. Traditionally, high-performance control has relied on methods like dynamical decoupling or filter-function engineering, which require at least partial characterization of the noise environment—a persistent experimental bottleneck. Our framework circumvents this requirement, decoupling control design from the arduous task of environmental characterization. This not only reduces experimental overhead but also provides resilience against uncharacterized or time-varying noise sources, a common challenge in real-world quantum devices.
@@
Looking forward, several avenues for future research are apparent. The current framework is derived under the Born-Markov approximation. Extending this approach to combat non-Markovian noise, which possesses memory effects, is a crucial next step and would broaden its applicability to an even wider range of physical systems. While our method relies on a first-order expansion of the noise effects, its formulation allows for the inclusion of higher-order terms from Eq. (10). Investigating the trade-off between the enhanced robustness from higher-order corrections and the increased computational cost of optimization would be a valuable pursuit.
Furthermore, the scalability of the optimization process presents a practical consideration. Although we have demonstrated success for a two-qubit gate, the computational resources required to calculate and minimize $D_{\rm eff}$ will grow with system size. Developing more efficient numerical techniques or leveraging machine learning to navigate the vast parameter space of many-qubit systems will be essential for applying this framework to larger-scale quantum processors.
@@
In conclusion, our universally robust control protocol provides a powerful and practical tool for mitigating decoherence in quantum technologies. Its hardware-agnostic nature and independence from noise-model specification offer a versatile solution applicable to quantum computing Berberich and Fink (2024), high-precision quantum sensing Hecht et al. (2025), and other quantum control tasks plagued by poorly characterized noise. By bridging a critical gap between theoretical control design and experimental reality, this work establishes a concrete and promising pathway toward the realization of fault-tolerant quantum information processing.

### v2

@@ II.4 Performance Evaluation of Universally Robust Control
To validate our framework, we implement the universally robust control protocol for two fundamental quantum tasks: (i) State transfer and (ii) Quantum gate operations under environmental noise. In the simulations of these control tasks, we model the environment as a thermal bath of harmonic oscillators with $\hat{H}_{\rm B}=\sum_{k}\left(\nu_{k}\hat{a}_{k}^{\dagger}\hat{a}_{k}+\frac{1}{2}\right)$, where $\hat{a}_{k}$ are bosonic annihilation operators for mode $k$ Breuer and Petruccione (2002). The system-environment coupling has the generic form $\hat{H}_{\rm I}=g\hat{A}\otimes\hat{B}$, where $\hat{B}=\sum_{k}\left(g_{k}/g\right)(\hat{a}_{k}+\hat{a}_{k}^{\dagger})$ and a super-Ohmic spectral density $J(\omega)=\omega_{\rm c}^{-2}\omega^{3}{\rm e}^{-\omega/\omega_{\rm c}}$ is assumed Weiss (2021). Here, $\omega_{\rm c}$ is the cutoff frequency, chosen as 10 times of the typical Bohr frequency of the bare system. Such a parameter choice corresponds to a large class of quantum systems Alkauskas et al. (2014); Jahnke et al. (2015); Norambuena et al. (2016).
Figure 1: Universally robust quantum state transfer in a two-level system. Quantum state transfer from initial state $\hat{\rho}_{-x}$ to target state $\hat{\rho}_{+x}$ under environmental noise. Dash-dotted lines: target-only control ($c=0$); solid lines: universally robust control ($c=10^{-2}$). (a) State infidelity $1-\mathcal{F}_{\rm state}$ versus coupling strength $\lambda$ for specific noise ($\hat{A}=\hat{\sigma}_{z}$). Inset: Infidelity versus $c$ for $\lambda=0.1$. (b) $1-\mathcal{F}_{\rm state}$ for generic noise ($\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$; $\mathbf{n}$: random unit vector). Data averaged over 100 realizations and the error bars indicate the standard deviation. (c) Optimized control fields $u_{x}(t)$ and $u_{y}(t)$ for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. The values of the time-integrated control power, $\bar{u}_{x,y}=\int_{0}^{\tau}u_{x,y}^{2}(t){\rm d}t$, are indicated in the legends. (d) State evolution trajectories on Bloch sphere for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. The evolution trajectory is color-coded to represent the purity of the state. Initial state: black square. Final states: blue dot (target-only) and green diamond (universally robust control). Parameters: Total control time $\tau=2/\Delta$ and inverse temperature $\beta=1/\Delta$.
@@ II.4.1 State Transfer in a Two-Level System
where $\hat{\sigma}_{i}$ ($i=x,y,z$) denote Pauli matrices and $u_{x,\,y}(t)$ correspond to independent control fields. Hereafter, we set the detuning $\Delta$ as the energy unit. This system exhibits complete controllability as the drift Hamiltonian ($\hat{\sigma}_{z}$) and the control Hamiltonians ($\hat{\sigma}_{x,\,y}$) generate the full $\mathfrak{su}(2)$ Lie algebra Huang et al. (1983); D’Alessandro (2022). Note that our chosen model with $\Delta\neq 0$ is the standard and most experimentally relevant one Chen et al. (2016); Kuzmanović et al. (2024).
@@
We implement quantum state transfer between orthogonal maximally coherent states: initial state $\hat{\rho}_{-x}=\frac{1}{2}(\hat{\mathbb{I}}-\hat{\sigma}_{x})$ and target state $\hat{\rho}_{+x}=\frac{1}{2}(\hat{\mathbb{I}}+\hat{\sigma}_{x})$. To evaluate protocol robustness, we consider two distinct noise channels: (i) specific coupling $\hat{A}=\hat{\sigma}_{z}$ and (ii) generic stochastic coupling $\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$ with $\mathbf{n}$ being a randomly oriented unit vector. Control fields were optimized by minimizing the objective functional $\mathcal{J}$ [Eq. (17)] via the CRAB algorithm, comparing target-only ($c=0$) versus robust ($c>0$) control strategies.
Fig. 1 highlights the striking advantages of our universally robust protocol. For a specific noise channel $\hat{A}=\hat{\sigma}_{z}$, while the infidelities for both protocols present an increasing tendency as $\lambda$ increases, their performance in completing the control task is distinctly different. Under increasing $\lambda$, the target-only optimization exhibits rapid fidelity degradation ($1-\mathcal{F}_{\text{state}}>0.3$ at $\lambda=0.1$), while the robust protocol maintains $1-\mathcal{F}_{\text{state}}<0.01$, achieving an infidelity two orders of magnitude lower than the target-only case, throughout $\lambda\in[10^{-4},10^{-1}]$. Note that in the true $\lambda=0$ limit, the target-only control ($c=0$) can yield a fidelity equal to or better than the robust control ($c>0$), as it optimizes only for noiseless infidelity $\mathcal{J}_{0}$. Besides, while a finite $c$ should definitely improve the robustness of the control against noise, it may still damage the final fidelity at some relatively large values due to the subtle trade-off between fidelity and robustness [see Eq. (17)]. This trade-off is clearly demonstrated in the inset of Fig. 1 (a), which reveals a broad “sweet spot” for $c$ (roughly in the range $10^{-4}$ to $10^{-1}$) where one can achieve a massive reduction in noise sensitivity.
@@
When faced with generic, uncharacterized noise $\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$ [Fig. 1 (b)], the universally robust protocol maintains consistently high fidelity across all 100 random realizations ($1-\mathcal{F}_{\text{state}}<0.02$ at $\lambda=0.1$), whereas the target-only approach fails catastrophically ($1-\mathcal{F}_{\text{state}}>0.3$). This clearly confirms the universality of our robust control protocol. Beyond its superior fidelity, the robust protocol is also significantly more efficient, reducing the required control field amplitudes by more than half compared to the target-only method [Fig. 1 (c)]. This efficiency is also reflected in the time-integrated control power (fluence), $\bar{u}_{x,y}=\int_{0}^{\tau}u_{x,y}^{2}(t){\rm d}t$, which is significantly lower for the robust protocol as shown in the figure’s legend. Furthermore, it excels at preserving quantum state purity. As illustrated by the Bloch sphere trajectories which are color-coded by state purity [Fig. 1 (d)], the robust evolution remains confined to the sphere’s surface, indicating a pure, unitary process, whereas the target-only path spirals inward—a clear signature of decoherence. These combined features demonstrate a capacity for universal noise resistance with minimal operational overhead.
This beneficial effect of achieving lower-amplitude and lower-fluence controls, as seen in Fig. 1 (c), is a direct physical consequence of our optimization. Our total objective function $\mathcal{J}$ [Eq. (17)] penalizes the noise sensitivity $D_{\rm eff}$. This metric $D_{\rm eff}$ is constructed from the kinetic coefficients $\gamma_{\alpha\alpha}[\omega_{j}(t)]$, which are related to the bath’s spectral density $J(\omega)$ evaluated at the system’s instantaneous Bohr frequencies $\omega_{j}(t)$. In our super-Ohmic model $J(\omega)\propto\omega^{3}{\rm e}^{-\omega/\omega_{\rm c}}$], this penalty is highly sensitive to large $\omega_{j}(t)$ values caused by “aggressive” pulses (large amplitudes and/or rapid oscillations). Consequently, by minimizing $D_{\rm eff}$, our optimizer is explicitly and strongly penalized for using such aggressive controls. It is therefore guided to find a “smoother” dynamical path to the target, which naturally corresponds to the gentler, lower-amplitude, and lower-fluence control fields observed in our results.
@@
Figure 2: Universally robust control for single-qubit Hadamard gate. Dash-dotted lines: target-only control ($c=0$); Solid lines: universally robust control ($c=10^{-2}$). (a) Gate infidelity $1-\mathcal{F}_{\rm gate}$ versus coupling strength $\lambda$ for specific noise ($\hat{A}=\hat{\sigma}_{z}$). Inset: Infidelity versus $c$ for $\lambda=0.1$. (b) $1-\mathcal{F}_{\rm gate}$ for generic noise ($\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$; $\mathbf{n}$: random unit vector). Data averaged over 100 realizations and the error bars indicate the standard deviation. (c) Optimized control fields $u_{x}(t)$ and $u_{y}(t)$ for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. The values of the time-integrated control power, $\bar{u}_{x,y}=\int_{0}^{\tau}u_{x,y}^{2}(t){\rm d}t$, are indicated in the legends. (d) State evolution trajectories on Bloch sphere under the Hadamard gate (from the $x$ to the $z$ direction) for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. The evolution trajectory is color-coded to represent the purity of the state. Initial state: black square. Final states: blue dot (target-only) and green diamond (universally robust control). Parameters: Total control time $\tau=2/\Delta$ and inverse temperature $\beta=1/\Delta$.
@@ II.4.2 Quantum Gate Operations
We extend our robust control protocol to the more demanding task of quantum gate synthesis, which requires noise-resilient state transformations across the entire Hilbert space—a significantly more complex challenge than single-state transfer Palao and Kosloff (2003); Aspman et al. . We demonstrate protocol effectiveness for two fundamental gates: the single-qubit Hadamard gate and the two-qubit controlled-Z (CZ) gate, which form a universal gate set when combined with T gates Nielsen and Chuang (2010).
@@
Using the control Hamiltonian from Eq. (18), we observe performance trends mirroring the two-level state transfer results. The influence of the noise strength $\lambda$ on the gate fidelity with $\hat{A}=\hat{\sigma}_{z}$ and $\hat{A}={\bf n}\cdot\hat{\boldsymbol{\sigma}}$ are depicted in Fig. 2 (a) and (b), respectively. In both cases, the target-only optimization results deviate substantially from the ideal value once $\lambda\neq 0$, while those of robust control exhibit strong noise robustness (e.g., preserving $1-\mathcal{F}_{\text{gate}}<0.02$ at $\lambda=0.1$ [Fig. 2 (b)]). Compared to the target-only optimization, the fidelity degradation of the robust protocol is suppressed by two orders of magnitude. The comparison of the required control fields in the two protocols is shown in Fig. 2 (c), where we find again that the robust optimization requires much lower control-field amplitudes and fluences than the target-only one. Besides, robust trajectories maintain confinement to the Bloch sphere surface (purity preservation), contrasting with target-only penetration (purity loss, as visually confirmed by the purity-coding of the trajectory) [Fig. 2 (d)].
Figure 3: Universally robust control for two-qubit CZ gate. Dash-dotted lines: target-only control ($c=0$); solid lines: universally robust control ($c=10^{-2}$). (a) Gate infidelity $1-\mathcal{F}_{\rm gate}$ versus coupling strength $\lambda$ for specific noise ($\hat{A}=\hat{\sigma}_{y}\otimes\hat{\sigma}_{0}$). Inset: Infidelity versus $c$ for $\lambda=0.1$. (b) $1-\mathcal{F}_{\rm gate}$ versus $\lambda$ for generic noise $\hat{A}=\sum_{\mu\nu}a_{\mu\nu}\hat{\sigma}_{\mu}\otimes\hat{\sigma}_{\nu}$, where the random coefficients $a_{\mu\nu}\sim N(0,1)$ with $\sum_{\mu\nu}a^{2}_{\mu\nu}=1$]. Data averaged over 100 realizations and the error bars indicate the standard deviation. (c) Optimized control fields $u_{j}(t)$ ($j=1,2,3,4$) for $\hat{A}=\hat{\sigma}_{y}\otimes\hat{\sigma}_{0}$ and $\lambda=0.1$. The values of the time-integrated control power, $\bar{u}_{j}=\int_{0}^{\tau}u_{j}^{2}(t){\rm d}t$, are indicated in the legends. Parameters: Total control time $\tau=2.5/\Delta$ and inverse temperature $\beta=1/\Delta$.
@@
where $\hat{\sigma}_{0}$ is the identity operator for a qubit. The control fields to be optimized are represented by $u_{j}(t)$ ($j=1,2,3,4$). Fig. 3 shows the variation of the gate fidelities against $\lambda$ for (a) $\hat{A}=\hat{\sigma}_{y}\otimes\hat{\sigma}_{0}$ and (b) $\hat{A}=\sum_{\mu\nu}a_{\mu\nu}\hat{\sigma}_{\mu}\otimes\hat{\sigma}_{\nu}$. Here, $\mu,\,\nu=0,x,y,z$, and the random coefficients $a_{\mu\nu}\sim N(0,1)$ with $\sum_{\mu\nu}a^{2}_{\mu\nu}=1$. In both cases, we find qualitatively the same behaviors as those of the single-qubit gate. That is, the infidelities of the target-only optimization increase rapidly as $\lambda$ increases, while the robust control can achieve and maintain high fidelity across a considerable range of noise amplitudes (e.g., preserving $1-\mathcal{F}_{\text{gate}}<0.01$ at $\lambda=0.1$ [Fig. 3 (b)]). Besides, as shown in Fig. 3 (c), the robust control requires significantly lowered control-field amplitudes and fluences. This efficiency, also seen in the lower time-integrated power, is a key advantage for scaling to multi-qubit systems where control power is a significant resource. The robust protocol consistently outperforms target-only optimization across all metrics, demonstrating scalability to multi-qubit systems while maintaining noise resilience and control efficiency.
@@ III Discussion
We have developed and validated a universally robust control framework capable of suppressing noise in open quantum systems without requiring specific knowledge of the noise channels. By engineering a control-objective functional that minimizes an intrinsic noise-susceptibility metric, our method achieves robustness against arbitrary Markovian noise within the first-order weak coupling approximation. Numerical simulations of state transfer and quantum gate operations show that this approach yields near-unity fidelities and suppresses errors by orders of magnitude compared to target-only optimization, even when subjected to diverse and poorly characterized noise sources.
[…]
@@
Looking forward, several avenues for future research are apparent. The current framework is derived under the Born-Markov approximation. Extending this approach to combat non-Markovian noise, which possesses memory effects, is a crucial next step and would broaden its applicability to an even wider range of physical systems. While our method relies on a first-order expansion of the noise effects, its formulation allows for the inclusion of higher-order terms from Eq. (10). Investigating the trade-off between the enhanced robustness from higher-order corrections and the increased computational cost of optimization would be a valuable pursuit. Regarding the aspect of noise sources, apart from the environmental noise considered in this paper, practical experimental platforms may also suffer from coherent errors associated with the unitary part of the dynamics. The origin of these errors can be traced back to system miscalibration Sheldon et al. (2016), sample aging Yan et al. , or the slow drift of an experimental setup Proctor et al. (2020). While a full analysis of these effects is beyond the scope of the current paper, our method may offer some incidental benefits. As we have found in Figs. 1-3 (c), the robust control solutions are characterized by lower amplitudes and smoother profiles, which are often less susceptible to hardware distortions and bandwidth limitations. We therefore hypothesize that our protocol may also exhibit enhanced resilience to certain types of coherent errors, and we frame this as a promising avenue for future investigation.
@@
Furthermore, the scalability of the optimization process presents a practical consideration. Although we have demonstrated success for a two-qubit gate, the computational resources required to calculate and minimize $D_{\rm eff}$ will grow with system size. Developing more efficient numerical techniques or leveraging machine learning to navigate the vast parameter space of many-qubit systems will be essential for applying this framework to larger-scale quantum processors. In conclusion, our universally robust control protocol provides a powerful and practical tool for mitigating decoherence in quantum technologies. Its versatile nature and independence from noise-model specification offer a flexible solution applicable to quantum computing Berberich and Fink (2024), high-precision quantum sensing Hecht et al. (2025), and other quantum control tasks plagued by poorly characterized noise. By bridging a critical gap between theoretical control design and experimental reality, this work establishes a concrete and promising pathway toward the realization of high-fidelity quantum information processing.
IV Methods
IV.1 Master equation and kinetic coefficients
To numerically implement the control protocol, it is convenient to rewrite the master equation Eq. (2) in a more concrete form which explicitly manifests the bath properties. For the quantum tasks we considered in the Results section, the system-bath interaction assumes the form $\hat{H}_{\rm I}=g\hat{A}\otimes\hat{B}$, and therefore the sum over the subscript $\alpha$ in Eq. (5) disappears, yielding
$\frac{{\rm d}}{{\rm d}t}\hat{\rho}_{\rm S}(t)=$ $-{\rm i}\left[\hat{H}_{\rm S}(t),\hat{\rho}_{\rm S}(t)\right]$ (22)
$+g^{2}\sum_{j}\left[\eta_{j}(t)\right]^{2}\gamma\left[\omega_{j}(t)\right]\left[\hat{F}_{j}(t)\hat{\rho}_{\rm S}(t)\hat{F}_{j}^{\dagger}(t)-\frac{1}{2}\{\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t),\hat{\rho}_{\rm S}(t)\}\right]\,.$
Substituting the bath operators $\hat{H}_{\mathrm{B}}=\sum_{k}\left(\nu_{k}\hat{a}_{k}^{\dagger}\hat{a}_{k}+\frac{1}{2}\right)$ and $\hat{B}=\sum_{k}\left(g_{k}/g\right)(\hat{a}_{k}+\hat{a}_{k}^{\dagger})$ into the kinetic coefficients, we have [see details below Eq. (5)]
$\gamma(\omega_{j})$ $=$ $\int_{-\infty}^{\infty}\mathrm{d}s\,\mathrm{e}^{-\mathrm{i}\omega_{j}s}\,\mathrm{Tr}_{\mathrm{B}}\left[\tilde{B}(s)\tilde{B}(0)\hat{\rho}_{\rm B}\right]$ (23)
$=$ $\int_{-\infty}^{\infty}\mathrm{d}s\,\mathrm{e}^{-\mathrm{i}\omega_{j}s}\,\sum_{k}\frac{g_{k}^{2}}{g^{2}}[(\overline{N}\left(\nu_{k}\right)+1)\mathrm{e}^{-\mathrm{i}\nu_{k}s}+\overline{N}\left(\nu_{k}\right)\mathrm{e}^{\mathrm{i}\nu_{k}s}]$
$=$ $\sum_{k}\frac{g_{k}^{2}}{g^{2}}[2\pi(\overline{N}\left(\nu_{k}\right)+1)\delta(\omega_{j}+\nu_{k})+2\pi\overline{N}\left(\nu_{k}\right)\delta(\omega_{j}-\nu_{k})]\,$,
[…]
$\gamma(\omega_{j})=2\pi\int_{0}^{\infty}\mathrm{d}\tilde{\omega}J(\tilde{\omega})[(\overline{N}\left(\tilde{\omega}\right)+1)\delta(\tilde{\omega}+\omega_{j})+\overline{N}\left(\tilde{\omega}\right)\delta(\tilde{\omega}-\omega_{j})].$ (24)
The rates for absorption $(\omega_{j}>0)$ and emission $(\omega_{j}<0)$ can be easily read from Eq. (24) as $\gamma(\left|\omega_{j}\right|)=2\pi J(\omega_{j})\overline{N}\left(\omega_{j}\right)$ and $\gamma(-\left|\omega_{j}\right|)=2\pi J(\left|\omega_{j}\right|)(\overline{N}\left(\left|\omega_{j}\right|\right)+1)$, respectively.
IV.2 Calculation of the Noise Sensitivity Metric $D_{\mathrm{eff}}$
We now briefly discuss how to obtain the noise sensitivity metric $D_{\mathrm{eff}}$ in practical calculations. The procedure is basically constituted by the following two steps:
1). The construction of the dynamical basis. The calculation begins with the numerical solution of the time-dependent Schrödinger equation for the system under a given control Hamiltonian $\hat{H}_{\rm S}(t)$. This yields the system propagator $\hat{U}_{\rm S}(t)$. At each time step of the numerical integration, $\hat{U}_{\rm S}(t)$ is diagonalized to determine the time-dependent eigenoperators (Lindblad jump operators) $\hat{F}_{j}(t)$ and the instantaneous Bohr frequencies $\omega_{j}(t)$. These form the natural, time-dependent basis in which the dissipative dynamics are described.
2). Numerical Integration and Norm Calculation. With the kinetic coefficients and constructed dynamical eigenoperators $\hat{F}_{j}(t)$, the operator $\mathbf{F}_{\alpha}$ in Eq. (14) is calculated by numerically integrating the term $\sum_{j}\gamma_{\alpha\alpha}[\omega_{j}(t_{1})]\tilde{\mathbf{F}}^{j}(t_{1})$ over the total evolution time $\tau$. Here, $\tilde{\mathbf{F}}^{j}(t)$ is the vectorized jump superoperator in the interaction picture. Finally, the value of $D_{\text{eff}}$ is computed from Eq. (14).
The above two steps serve as a practical recipe for implementing our method.
DATA AVAILABILITY
Data is available from the corresponding author upon reasonable request.
CODE AVAILABILITY
The codes used to generate data for this paper are available from the corresponding author upon reasonable request.